Bukit Barisan (04.01.2014 to 06.30.2014)

You are here

PDF versionPDF version
Year: 
2014
Quarter: 
April - June
Communications activities: 

 No news

Unusual Events at the TEAM Site: 

 No news

New Species: 

 No news

Protocol Activities: 

 

Terrestrial Vertebrate Protocol: The camera traps data was collected in 3 arrays (the latest pick up in July 2014), currently the photos are still being identified. 

New camera traps without ID. (need to get an update from TEAM for the ID, the serial numbe: H500 EE 05128009 and H500 EE 05128389

Camera traps with error signs:RM13AC02003087, RM13AC02003098, RM13AC02003060

Missing: RM13AC02003080 

Considered the current results, all cameras and their accessories need replacement.

Vegetation protocol: During this quarter the vegetation monitoring team has measured for three plots VGBBS 1,3 and 5 (completed in July 2014). The other three plots will be collected on August 2014.

Climate Protocol: since June 2013 climate station is not operational, damaged by Elephant and no replacement to date. This had been reported in report period 10.01.2013-12.31.2013.

 

Protocol Problems: 

There were limited number of camera traps available so the last camera trap surveys were conducted in 3 arrays, and cover all points. We are currently in dicussion on how to upload the data as the DeskTEAM only has 2 arrays.

Two new cameras were also used, but still with no Camera ID. 

Schedule Problems: 

 There was a delayed in collecting data for vegetation protocol, simply because in this period we were trying to cover all point for the camera traps hence the camera trap surveys were conducted in three array. The previous site manager had been communicated this with TEAM.